# PriorRG: Prior-Guided Contrastive Pre-training and Coarse-to-Fine Decoding for Chest X-ray Report Generation

### v1

@@ B.3 Qualitative Analysis
Additional qualitative analysis is presented in Figure 7. The results show that PriorRG outperforms the baseline (Liu et al. 2024b) in generating context-aware, disease progression-sensitive radiology reports. In cases with the prior image but no clinical context (Case 1), the model effectively captures disease progression by describing pleural effusions as stable over time. When only clinical context is available—such as chest pain radiating to the back (Case 2) or hypoxia (Case 4)—PriorRG generates radiology reports that align with the underlying clinical intent, accurately reflecting surgical history (e.g., CABG in Case 2) and comprehensively assessing relevant cardiopulmonary structures in response to hypoxia (Case 4). Even in the absence of both the prior image and clinical context (Case 3), our PriorRG still produces complete and clinically plausible reports.
Figure 7: Qualitative comparison of reports from the baseline (Liu et al. 2024b), our proposed PriorRG, and radiologists. Sentences in the radiologist-generated report are color-coded to match corresponding descriptions in the AI-generated reports. In PriorRG’s output, correct disease progression patterns and failure descriptions are marked in bold and underlined, respectively.
B.4 Failure Case Analysis
@@
Figure 7 illustrates that our PriorRG occasionally produces overstatements, such as simplifying mildly enlarged cardiac silhouette to enlarged (Case 3). This phenomenon mainly stems from the model’s lack of effective modeling of fine-grained medical attributes, leading to insufficient control over the severity and semantic accuracy in generated reports. To address this limitation, we are exploring attributed abnormality graphs (Zhang et al. 2024; Yan et al. 2023) to more accurately capture and represent attribute-specific disease states.

### v2

@@ B.4 Qualitative Analysis
Additional qualitative analysis is shown in Figure 7. The results show that PriorRG outperforms the baseline (sei) in generating context-aware, disease progression-sensitive radiology reports. In cases with the prior image but no clinical context (Case 1), the model effectively captures disease progression by describing pleural effusions as stable over time. When only clinical context is available—such as chest pain radiating to the back (Case 2) or hypoxia (Case 4)—PriorRG generates radiology reports that align with the underlying clinical intent, accurately reflecting surgical history (e.g., CABG in Case 2) and comprehensively assessing relevant cardiopulmonary structures in response to hypoxia (Case 4). Even in the absence of both the prior image and clinical context (Case 3), our PriorRG still produces complete and clinically plausible reports.
B.5 Failure Case Analysis
@@
Figure 7 illustrates that our PriorRG occasionally produces overstatements, such as simplifying mildly enlarged cardiac silhouette to enlarged (Case 3). This phenomenon mainly stems from the model’s lack of effective modeling of fine-grained medical attributes, leading to insufficient control over the severity and semantic accuracy in generated reports. To address this limitation, we are exploring attributed abnormality graphs (tmi-2024-scene-graph-rrg; yan_tmi_2023) to more accurately capture and represent attribute-specific disease states.